OpenStudy (rebeccapink6398):

Can someone please help? <3 60% of a number is 90. What is 50% of that same number?

Vocaloid (vocaloid):

60% is equal to 0.6 0.6*x = 90 solve for x, then find 50% of x

OpenStudy (kyle.ingoldsby415):

60%=60/100=0.6 0.6*x=90 x=90/0.6 x=150 -> 60% of 150 = 0.6*150 = 90

OpenStudy (anonymous):

its 75

OpenStudy (rebeccapink6398):

YAY!!! Thanks @kyle.ingoldsby415 and thanks @xavierbo2

OpenStudy (anonymous):

90 divided by 60/100 which is 90/0.6 which is 15

OpenStudy (kyle.ingoldsby415):

ur welcome

OpenStudy (rebeccapink6398):

65% of what number is 39?

OpenStudy (kyle.ingoldsby415):

39 is 65% of what number 65%= 0.65 39/0.65= 60

OpenStudy (rebeccapink6398):

so the answer is 60?... oh it make ssence lol i was multiplying

OpenStudy (mathstudent55):

You can also use a proportion. 60% is 0.60 as a decimal 50% is 0.50 as a decimal the unknown number is x the 50% of the number we are looking for is y 0.6x is to 90 as 0.5x is to y \(\dfrac{0.6x}{90} = \dfrac{0.5x}{y} \) Divide both sides by x, and the only unknown left is y: \(\dfrac{0.6}{90} = \dfrac{0.5}{y} \) Cross multiply: \(0.6y = 0.5 \times 90\) \(0.6y = 45\) \(y = \dfrac{45}{0.6} \) \(y = 75\) Answer: the number is 75